# Test-time Adaptive Hierarchical Co-enhanced Denoising Network for Reliable Multimodal Classification

### v1

@@ IV-D3 Further Discussion On the Subspace Projection Alignment
IV-E Discussion On SACA
IV-E1 Effectiveness of Noise Experts
More experiments are conducted to verify that the modality specific noise experts and cross modality noise experts in SACA respectively facilitate the removal of instance level modality-specific and cross-modality noise. Specifically, we first introduce only modality-specific noise $\epsilon=5$ into one modality of the dataset and visualize the sample masks $w_{s}^{m}$ learned by the corresponding modality-specific noise expert. Fig. 5 illustrates the masks learned by the modality-specific noise expert on the image modality affected by noise in the CUB dataset. In the noise-free setting (left), the mask values exhibit strong consistency across samples. After modality-specific noise is added (right), the learned mask show increased variability between samples, reflecting the instance-level diversity of the modality-specific noise and demonstrating that the modality-specific noise expert effectively captures the noise for each sample. Also, we introduce cross-modality noise with $\eta=20\%$ to one modality of the dataset by randomly shuffling the features of 20% of its samples. Fig. 6 shows the masks learned by the cross-modality noise expert on the image modality affected by noise in the CUB dataset. Before noise is added (left), the masks of different samples exhibit a relatively consistent pattern, whereas after noise introduction (right), the masks show substantial variability across samples. This demonstrates that the cross-modality noise expert can effectively capture sample-specific corrupted information introduced by noise.
IV-E2 Discussion on Confidence-Aware Asymmetric Slack Alignment
We first conduct experiments to validate the effectiveness of the slack alignment proposed in Confidence-Aware Asymmetric Slack Alignment. Specifically, we compare with existing methods that enforce cross-modality consistency, including similarity-based approaches (sim), as used in [22, 18, 43], and information-theoretic approaches (MI), as used in [44]. “slack” represents applying our $\mathcal{L}_{\mathrm{nll}}^{c}$ that retains the slack-alignment mechanism while removing modality-confidence guidance. As shown in Table VIII, under cross-modality noise with varying intensity $\eta$, our slack alignment exhibits stronger stability than other methods, indicating superior robustness. “MI” achieves better performance than “sim”, likely due to its less restrictive constraint on cross-modality consistency.
We also conduct experiment to validate the effectiveness of incorporating modality confidence to guide slack alignment. Specifically, the model is trained on data with modality-specific and cross-modality noise ($\epsilon=1,\eta=10\%$), and the changes in modality-wise confidence and the loss $\mathcal{L}_{\mathrm{nll}}^{c}$ are reported for both with and without confidence guidance. The variant without confidence guidance is implemented by removing the gradient modulation in Eq. (13). Modality confidence reported in Fig. 7 is computed as the average of the likelihoods $p^{m}(\hat{h}_{s,i}^{m})$ of all sample representations $\hat{h}_{s,i}^{m}$ under the corresponding modality-specific feature prior $p^{m}$. As shown in the left panel of Fig. 7, as training progresses, the confidence of each modality gradually increases, indicating progressive noise removal. Moreover, with modality-confidence guidance, modality confidence rises to higher levels than without such guidance, suggesting that confidence guidance promotes alignment toward high-confidence features across modalities, thereby yielding more robust representations. Moreover, as shown in the right panel of Fig. 7, the loss $\mathcal{L}_{\mathrm{nll}}^{c}$ decreases faster with confidence guidance than without it, indicating that confidence guidance also accelerates the alignment process.
Figure 7: Changes in modality-wise confidence and the decline of $\mathcal{L}_{\mathrm{nll}}^{c}$ during training on the FOOD101 dataset with cross-modality noise $(\eta=10\%)$ and modality-specific noise $(\epsilon=1)$, under two conditions: with and without the guidance of modality confidence on slack alignment.
TABLE VIII: Comparison of model performance under noise using different cross-modality alignment methods.
Dataset Method $\eta=0$ $\eta=10\%$ $\eta=20\%$
BRCA sim 84.6 66.8 57.3
MI 85.0 70.8 61.3
slack 85.6 73.4 65.8
ROSMAP sim 86.2 75.3 62.8
MI 86.9 76.4 65.1
slack 87.7 79.2 71.8
CUB sim 92.0 69.8 60.5
MI 92.3 70.6 62.7
slack 92.8 73.0 66.2
FOOD101 sim 92.5 71.3 61.8
MI 93.0 72.6 64.1
slack 93.2 75.5 68.4
IV-F Discussion On TTCE
@@
To verify that TTCE effectively enhances the model’s adaptability and generalization to previously unseen noise, we conduct experiments using the original training data and test data corrupted with modality-specific and cross-modality noise ($\eta=10\%,\epsilon=5$). The model is first trained, and then the representations $\hat{h}_{s}^{m}+\hat{h}_{c}^{m}$ of each modality during the iterative enhancement of TTCE on the noisy test data are visualized. As shown in Fig. 4, before TTCE iterations (0 iteration), the learned representations exhibit poor intra-modality class separability and weak cross-modality alignment (high $\mathcal{L}_{\mathrm{nll}}^{c}$), indicating insufficient removal of newly encountered noise. As TTCE proceeds, intra-modality class separability gradually improves, and cross-modality alignment is substantially enhanced ($\mathcal{L}_{\mathrm{nll}}^{c}$ decreases), indicating that newly introduced modality-specific and cross-modality noise is progressively removed. These results demonstrate that, through multiple TTCE updates, the model can progressively adapt to and remove noise unseen in training stage, thereby achieving stronger generalization performance.
Figure 8: Variation of test accuracy under different values of $\alpha$ and $\beta$.
@@
IV-G Parameter Analysis
More experiments are conducted to analyze the hyperparameters of the proposed model. In these experiments, TAHCD is trained on the original dataset and test on data with modality-specific noise ($\epsilon=5$) and cross-modality noise ($\eta=10\%$). TAHCD mainly involves three hyperparameters, namely the update ratios $\alpha$ and $\beta$ for $\mu$ and $\Sigma$ in Eq. (17), and the number of update iterations $E$ in TTCE. The values of $\alpha$ and $\beta$ lie in $[0,1]$, indicating the extent to which the modality feature distributions incorporate the newly estimated mean and covariance. Setting $\alpha=\beta=0$ means that the distributions are not updated, while $\alpha=\beta=1$ indicates that the newly computed statistics fully replace the original ones. As shown in Fig. 8, increasing $\alpha$ and $\beta$ leads to a model classification accuracy trend that first improves and then degrades. The $\alpha$ and $\beta$ values that achieve the best performance are relatively consistent across datasets, for example, $\alpha=0.4$ and $\beta=0.3$ on BRCA, and $\alpha=0.4$ and $\beta=0.4$ on ROSMAP. Additionally, Fig. 9 shows that as the number of update iterations $E$ increases, the test accuracy first rises and generally stabilizes around $E=30$ across all datasets. This indicates that the model has completed its adaptive adjustment to the noise present in the current test data. When the noise severity increases, more update iterations are required, and the optimal value of $E$ corresponding to the highest accuracy increases accordingly.
V Conclusion
@@
Existing multimodal learning methods struggle to achieve robust multimodal representation learning under heterogeneous noise and exhibit limited adaptability and generalization to unseen noise. To address these challenges, we propose the Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD). TAHCD introduces Adaptive Stable Subspace Alignment and Sample-Adaptive Confidence Alignment to jointly remove modality-specific and cross-modality noise at the global and instance levels, respectively, thereby improving robustness. In addition, TAHCD designs a test-time cooperative enhancement mechanism that enables colaborative enhancement between global-level and instance-level denoising processes driven by input sample noise in a label-free manner. This enables TAHCD to adapt more effectively to unseen noise at test time, improving the model’s generalization ability. Experiments on multiple benchmarks demonstrate that TAHCD outperforms other state-of-the-art reliable multimodal learning methods in robustness and generalization ability.
@@
Although existing reliable multimodal learning methods have been extensively studied for classification tasks, their reliability in other machine learning tasks, such as regression and semantic segmentation, remains insufficiently explored. This presents a promising and important direction for further investigation. Accordingly, future work will extend and validate the proposed approach across a broader set of machine learning tasks.

### v2

@@ 4.5.2 Discussion on Confidence-Aware Asymmetric Slack Alignment
4.6 Discussion On TTCE
4.6.1 Adaptation And Enhancement against Unseen Noise
@@
To verify that TTCE effectively enhances the model’s adaptability to previously unseen noise, we conduct experiments using the original training data and test data corrupted with modality-specific and cross-modality noise ($\eta=10\%,\epsilon=5$). The model is first trained, and then the representations $\hat{h}_{s}^{m}+\hat{h}_{c}^{m}$ of each modality during the iterative enhancement of TTCE on the noisy test data are visualized. As shown in Figure 7, before TTCE iterations (0 iteration), the learned representations exhibit poor intra-modality class separability and weak cross-modality alignment (high $\mathcal{L}_{\mathrm{nll}}^{c}$), indicating insufficient removal of newly encountered noise. As TTCE proceeds, intra-modality class separability gradually improves, and cross-modality alignment is substantially enhanced ($\mathcal{L}_{\mathrm{nll}}^{c}$ decreases), indicating that newly introduced modality-specific and cross-modality noise is progressively removed. These results demonstrate that, through multiple TTCE updates, the model can progressively adapt to and remove noise unseen in training data, thereby achieving stronger generalization performance.
@@
Table 7: Comparison of generalization performance and runtime across methods on the BRCA dataset. Cross-modality noise and modality-specific noise with severities $\eta$ and $\epsilon$ are added only onto the test data. The runtime for TAHCD is reported with the best TTCE iteration $E$ under each noise condition shown in parentheses.
Data Type Accuracy and Efficiency MD PDF SPS TAHCD
$\eta=0$ $\epsilon=0$ ACC (%) 82.1 83.7 79.8 85.6
Runtime (ms) 160.22 124.67 52.03 99.23 ($E=5$)
$\eta=5\%$ $\epsilon=1$ ACC (%) 69.6 72.1 66.3 79.5
Runtime (ms) 160.22 124.67 52.03 136.35 ($E=20$)
$\eta=10\%$ $\epsilon=5$ ACC (%) 51.4 56.6 46.1 67.4
Runtime (ms) 160.22 124.67 52.03 165.28 ($E=30$)
Table 8: Top 5 informative biomarkers that TAHCD learned from different modalities on BRCA dataset.
Modality Top 5 informative biomarkers
DNA ZNF671, KRTAP3-1, AGR2, MYT1, RPS6KL1
mRNA KLK8, PTX3, SOX11, KRT6B, RDH10
miRNA mir-21, mir-190b, mir-378, mir-93, mir-23b
4.6.2 Computational Efficiency
Additional experiments are conducted to evaluate the computational efficiency of TTCE under noisy conditions. Since TTCE does not introduce additional network structures or parameters, we focus on evaluating runtime of the model during testing. As shown in Table 7, modality-specific and cross-modality noise with varying severities $\epsilon$ and $\eta$ are added into the two modalities of the CUB testing data, and we compare the test accuracy and runtime of different methods. We evaluate the test accuracy of TAHCD every five enhancement iterations in the experiment. It can be observed that TAHCD achieves superior and more stable performance while maintaining competitive runtime, demonstrating its high computational efficiency. Under noise-free condition ($\eta=0,\epsilon=0$), TAHCD exhibits lower runtime than MD and PDF. As noise increases, TTCE in TAHCD requires more enhancement iterations $E$, resulting in longer runtime. Even so, under $\eta=10\%$ and $\epsilon=5$, TAHCD maintains runtime comparable to MD and PDF while achieving over 10% higher accuracy.
4.6.3 Discussion on the Number of Enhancement Iterations $E$
We add modality-specific noise with $\epsilon=5$ and cross-modality noise with $\eta=10\%$ into the test sets of four datasets and report the test accuracy of TAHCD under different values of $E$. Figure 8 shows that as the number of update iterations $E$ increases, the test accuracy first rises and generally stabilizes around $E=30$ across all datasets. This indicates that the model has completed its adaptive adjustment to the noise present in the current test data. Moreover, Table 7 shows that when the noise severity increases, the optimal value of $E$ corresponding to the highest test accuracy increases, indicating that more enhancement iterations are required.
4.7 Biomarkers Identification
To validate that TAHCD reliably learns informative features, following previous works [13, 43], we examine the biomarkers identified by TAHCD for early diagnosis and prognosis on the multi-omic BRCA dataset. Specifically, we compute each modality’s feature informativeness as the mean of the product between the global-level informativeness $w_{\lambda}^{m}$ learned in Eq. (5) and the instance-level informativeness $w_{s,i}^{m}$ learned in Eq. (10) by TAHCD. The results in Table 8 show that the biomarkers identified by TAHCD are supported by existing medical studies as key indicators for cancer diagnosis and prognosis. For instance, KLK8 has been reported as an independent prognostic indicator in breast cancer patients [23], while low expression of ZNF671 [35] and miR-378 [2] are both linked to poor prognosis in breast cancer.
Figure 9: Variation of test accuracy under different values of $\alpha$.
4.8 Parameter Analysis
More experiments are conducted to analyze the hyperparameter $\alpha$ of TAHCD defined in Eq. (17). $\alpha\in[0,1]$ indicates the extent to which the modality feature distributions incorporate the newly estimated mean and covariance, where $\alpha=0$ means that the distributions are not updated, and $\alpha=1$ indicates that the newly computed statistics fully replace the original ones. Figure 9 shows that as $\alpha$ increases, the accuracy of TAHCD first rises and then declines. The model achieves substantially better performance when $\alpha\neq 0$ than when $\alpha=0$, demonstrating the effectiveness of TAHCD’s adaptive updates in improving generalization to unseen noise. Moreover, TAHCD achieves high accuracy across a relatively wide range of $\alpha$, indicating low sensitivity to this hyperparameter.
5 Conclusion
@@
Existing multimodal learning methods struggle to achieve robust multimodal representation learning under heterogeneous noise and exhibit limited adaptability and generalization to unseen noise. To address these challenges, we propose the Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD). TAHCD introduces Adaptive Stable Subspace Alignment and Sample-Adaptive Confidence Alignment to jointly handle modality-specific and cross-modality noise at the global and instance levels, respectively, thereby improving robustness. In addition, TAHCD designs a Test-Time Cooperative Enhancement mechanism that enables cooperative enhancement between global-level and instance-level denoising processes driven by input sample noise in a label-free manner. This enables TAHCD to adapt more effectively to unseen noise at test time, improving the model’s generalization ability. Experiments on multiple benchmarks demonstrate that TAHCD outperforms other state-of-the-art reliable multimodal learning methods in robustness and generalization ability.
@@
Although existing reliable multimodal learning methods have been extensively studied for classification tasks, their reliability in other machine learning tasks, such as regression, remains insufficiently explored. This presents a promising and important direction for further investigation. Accordingly, future work will extend and validate the proposed approach across a broader set of machine learning tasks.